OpenStudy (anonymous):

30 year 4/1 ARM at 5.5% with a 2/6 cap structure Initial monthly payment: $1,987.26 What will be the balance of the loan at the end of the initial interest rate period?

OpenStudy (amistre64):

the same as the balance for a fixed 30 year loan after 4 years

OpenStudy (anonymous):

Marcos obtains a mortgage for $350,000 with the following terms:

OpenStudy (anonymous):

$329,484.18 $254,611.45 $323,612.11 $320,806.22

OpenStudy (amistre64):

\[\Large B_o(1+\frac{.055}{12})^{4*12}-P\frac{1-(1+\frac{.055}{12})^{4*12}}{1-(1+\frac{.055}{12})}\]

OpenStudy (anonymous):

these are the answers that i have to choose from

OpenStudy (anonymous):

can you help me

OpenStudy (amistre64):

i just did ... use that setup by plugging in the initial balance and the payments

OpenStudy (anonymous):

$323,612.11 thats what i got.. is it correct??

OpenStudy (amistre64):

thats not what i get .... im about 6000 more than that

OpenStudy (amistre64):

your results are for 5 years, not 4

OpenStudy (anonymous):

okay i think i see what i did wrong i got $329,484.18

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

youre welcome